1. Determine the dimension of and a basis for the solution space of the homogeneous linear system.

2x subscript 1 + x subscript 2 + 2x subscript 3 = 0
x subscript 1 + 6x subscript 3 = 0
x subscript 2 + x subscript 3 = 0

2. Find a basis for the given subspace of R Superscript 3, and state its dimension for the plane 2 x - 3 y + 5 z = 0.

3. Let v1 = (1,-1,3,-6) and v2 = (-3,4,-12,24).
Find standard basis vectors for R4 that can be added to the set {v1, v2} to produce a basis for R4.

A. v3 = (1,0,0,0) and v4 = (0,1,0,0)
B. v3 = (0,0,1,0) and v4 = (1,0,0,0)
C. v3 = (1,0,0,0) and v4 = (0,0,0,1)
D. v3 = (0,1,0,0) and v4 = (0,0,0,1)
1
Expert's answer
2020-01-12T15:03:17-0500

"2x_1+x_2+2x_3=0,"

"x_1+6x_3=0,"

"x_2+x_3=0."

"(R_1 \gets R_1-2R_2)"

"x_2+2x_3-12x_3=0,"

"x_1+6x_3=0,"

"x_2+x_3=0."

"(R_1 \gets R_1-R_3)"

"-11x_3=0,"

"x_1+6x_3=0,"

"x_2+x_3=0."

Thus, "x_1=0, x_2=0, x_3=0."

A basis for the solution space is "(0,0,0)," the dimension of this space in "R^3" is 0.

2."2x=3y-5z"

"x=3y\/2-5z\/2"

"\begin{pmatrix}
 ( 3/2) y-(5/2)z \\
 y\\
z
\end{pmatrix}=\begin{pmatrix}
 (3/2) \\
 1 \\
 0
\end{pmatrix}y+\begin{pmatrix}
 - 5/2 \\
 0\\
 1
\end{pmatrix}z"

{"v_1,v_2" }={(3/2,1,0),(5/2,0,1)} spans R3

We just have to prove that "a_1,a_2=0" in "a_1v_1+a_2v_2=0"

"a_1(3\/2,1,0)+a_2(-5\/2,0,1)=0"

"(3\/2)a_1-(5\/2)a_2=0"

"a_1=0"

"a_2=0"

So, { "v_1,v_2" } is the basis of the subspace in R3, the dimension of the subspace in "R^3" is 2.

3.(D)

"v_1=(1,-1,3,-6)"

"v_2=(-3,4,-12,24)"

"v_3=(x_1,y_1,z_1,w_1)"

"v_4=(x_2,y_2,z_2,w_2)"

"a_1v_1+a_2v_2+a_3v_3+a_4v_4=0"

"a_1 (1,-1,3,-6) +a_2(-3,4,-12,24)+a_3(x_1,y_1,z_1,w_1)+a_4 (x_2,y_2,z_2,w_2) =0"

From above we obtain the following equation:

"a_1-3a_2+x_1a_3+x_2a_4=0"

"-a_1+4a_2+y_1a_3+y_2a_4=0"

"3a_1-12a_2+z_1a_3+z_2a_4=0"

"-6a_1+24a_2+w_1a_3+w_2a_4=0"

From the options we put the values of v3 and v4 in the above equations.

"a_1=a_2=a_3=a_4=0" only for D option.